# Multiplayer Federated Learning: Reaching Equilibrium with Less Communication

### v1

@@ 3 PEARL-SGD: Algorithm and Convergence Guarantees
In this section, we introduce and analyze Algorithm 1, named Per-Player Local SGD (PEARL-SGD), which is suitable for the MpFL setting we described in Section 2.
@@ 3.1 Algorithm and assumptions
PEARL-SGD works by having the clients/players of the game run SGD independently in parallel for updating their strategy (keeping the strategies $x^{-i}$ of the other players fixed) and concatenate the strategies of all players only once in a while (via a central server). In more detail, in every round of PEARL-SGD, each player $i\in[n]$ runs $\tau$ iterations of stochastic gradient descent (SGD) with respect to $f_{i}(\cdot,x^{-i})$, having the $x^{-i}$ fixed to be the information of the other players’ actions obtained from the previous synchronization step. Once each player completes $\tau$ iterations of SGD (local updates), a synchronization occurs. In each synchronization, the central server collects the actions of all players, and then the concatenation of all updated strategies/actions is distributed to all clients/players.
@@
Input: Step-sizes $\gamma_{k}>0$, Synchronization interval $\tau\geq 1$, Number of synchronization/local update rounds $R\geq 1$
@@
     Master server collects $x_{\tau p}^{i}$ from players $i=1,\dots,n$ and forms ${\mathbf{x}}_{\tau p}=(x_{\tau p}^{1},\dots,x_{\tau p}^{n})$
     Master server distributes ${\mathbf{x}}_{\tau p}$ back to players $i=1,\dots,n$
@@
Let us note that the synchronization step in PEARL-SGD involves transferring a $D=(d_{1}+\dots+d_{n})$-dimensional vector ${\mathbf{x}}_{\tau p}$ from the master server to the players. This is an important difference compared to the classical FL (minimization problem), where the dimension of the communication vectors is the same from client to master and from master to client, and it does not scale with $n$. While PEARL-SGD aims to reduce this overhead compared to its distributed variant ($\tau=1$) by communicating less frequently (with $\tau>1$), the high complexity of the synchronization step makes MpFL (and distributed $n$-player games in general) more suitable for cross-silo FL setups with relatively small number of organizations and more reliable communication bandwidths. We expect that the potentially expensive communication of high-dimensional vectors ${\mathbf{x}}_{\tau p}$ could be addressed by incorporating additional techniques such as gradient compression [3, 8]. This is an orthogonal approach to our proposed local methods, and we leave it for future work.
@@
There exists a unique equilibrium ${\mathbf{x}}_{\star}=(x_{\star}^{1},\dots,x_{\star}^{n})\in\mathbb{R}^{D}$, for which ${\varbbF}({\mathbf{x}}_{\star})=0$, and $\mu>0$ such that for any ${\mathbf{x}}\in\mathbb{R}^{D}$, $\left\langle{\varbbF}({\mathbf{x}}),{\mathbf{x}}-{\mathbf{x}}_{\star}\right\rangle\geq\mu\left\|{\mathbf{x}}-{\mathbf{x}}_{\star}\right\|^{2}$.
(QSM) is a concept extending quasi-strong convexity [37, 36] to the context of variational inequality problems (VIPs). This condition has been referred to as different names in the literature, such as strong coherent VIPs [108], VIPs with strong stability condition [80], or the strong Minty variational inequality [22]. It generalizes strong monotonicity, capturing some non-monotone problems. In [75], it was proposed and utilized as an assumption ensuring the convergence of SGDA dynamics in minimax games without the well-known issues of cycling or diverging [81, 16]. Later, it was also used in the analysis of stochastic extragradient method [33] and its single-call variants (optimistic and past stochastic etragradient) [13].
@@
(SCO) generalizes the class of coercive operators and, interestingly, can hold for non-Lipschitz operators [75]. This has also been taken as minimal assumption for SGDA analysis in prior work [8]. Note that (QSM) and (SCO) together imply $\mu\left\|{\mathbf{x}}-{\mathbf{x}}_{\star}\right\|\leq\left\|{\varbbF}({\mathbf{x}})\right\|\leq\ell\left\|{\mathbf{x}}-{\mathbf{x}}_{\star}\right\|$ for any ${\mathbf{x}}\in\mathbb{R}^{D}$, which implies $\mu\leq\ell$. We call $\kappa=\ell/\mu\geq 1$ the condition number of the problem. In Appendix D, we provide a detailed discussion on the set of our theoretical assumptions and explain connections to other commonly assumed properties in the literature such as cocoercivity, Lipschitzness and monotonicity.
@@ 3.2 Convergence of PEARL-SGD: Deterministic setup
First, we provide the convergence result for PEARL-SGD with constant step-size $\gamma_{k}\equiv\gamma$ in the full-batch (deterministic) scenario, where there is no noise in the gradient computation. While this is recovered as a special case of Theorem 3.4, we state it separately because the deterministic case provides several points of discussion that are worth emphasizing on their own.
@@
Assume (CVX), (SM), (QSM) and (SCO). Let $L_{\mathrm{max}}=\max\{L_{1},\dots,L_{n}\}$, $0<\gamma_{k}\equiv\gamma\leq\frac{1}{\ell\tau+2(\tau-1)L_{\mathrm{max}}\sqrt{\kappa}}$, and let $\kappa=\ell/\mu$ be the condition number. Then the Deterministic PEARL-SGD (Algorithm 1 with full-batch) converges with the rate
@@
Theorem 3.3 shows that deterministic PEARL-SGD converges linearly to an equilibrium. This distinguishes our result from the analyses of local gradient descent for finite sum minimization in heterogeneous data setups, where one has convergence to a neighborhood of optimum even when there is no noise [51], unless further correction mechanism is used [84]. In addition, note that when $\tau=1$, the step-size constraint and the convergence rate of Theorem 3.3 coincide with those from the analysis of the gradient descent-ascent (GDA) under the (QSM) and (SCO) assumptions from [75], showing the tightness of our analysis.
@@
If $\gamma$ does not appropriately scale down with $\tau$, then at each round, players’ actions (SGD iterates) converge to minimizers of local functions. We call this phenomenon player drift, analogous to client drift in classical FL [50], enforcing the ${\mathcal{O}}(1/\tau)$ step-size. In our setting, note that the local minimizers $x^{i}_{\star}(x_{\tau p}^{-i}):=\operatorname*{argmin}_{x^{i}\in\mathbb{R}^{d_{i}}}f_{i}(x^{i};x_{\tau p}^{-i})$ depend on other players’ strategies $x_{\tau p}^{-i}$. Due to this dependence, under extreme player drift, PEARL-SGD may display undesirable dynamics such as diverging to infinity [^1] This can be checked with simple examples such as the two-player quadratic minimax game $\min_{u\in\mathbb{R}}\max_{v\in\mathbb{R}}\,\frac{\mu}{2}u^{2}+uv-\frac{\mu}{2}v^{2}$ with $\mu<1$.. As these features are not typically observed in client drift in classical FL, player drift represents a distinct phenomenon despite some conceptual similarities. Therefore, we consider understanding and mitigating player drift an intriguing direction for future work in MpFL, which may necessitate novel insights that differ from existing approaches to client drift [50, 84].
@@ 3.3 Convergence of PEARL-SGD: Stochastic setup
We now discuss the convergence of PEARL-SGD with stochastic gradients. We first present the convergence of PEARL-SGD to a neighborhood of an equilibrium ${\mathbf{x}}_{\star}$ given constant step-sizes $\gamma_{k}\equiv\gamma$, and then discuss the communication complexity gain we achieve. Then we present the convergence result using a decreasing step-size selection, showing sublinear convergence to the exact equilibrium ${\mathbf{x}}_{\star}$ rather than its neighborhood. While we defer the details of the proofs to Appendix B, we provide a proof outline for Theorem 3.4 in Section 3.4.
@@
If we use the largest possible step-size $\gamma=\frac{1}{\ell\tau+2(\tau-1)L_{\mathrm{max}}\sqrt{\kappa}}$ allowed in Theorem 3.4, then the right-hand side of the bound does not scale down indefinitely with $\tau$. In fact, with this choice of $\gamma$, one can expect the communication gain by a factor of approximately $\frac{L_{\mathrm{max}}}{\ell}$ (when $L_{\mathrm{max}}\ll\ell$). More precisely, suppose $q\leq 1$ (equivalently $L_{\mathrm{max}}\leq\sqrt{\ell\mu}$—refer to Appendix D for the explanation that this is a common parameter regime). Then we have $\gamma=\Theta\left(\frac{1}{\ell\tau}\right)$ and
@@
In Corollary 3.5, we analyze the convergence and communication gain of PEARL-SGD in the regime where the total number of iterations $T=\tau R$ is large. In this setting, we use a step-size that depends on the total number of iterations $T$.
@@
Suppose the assumptions of Theorem 3.4 hold, and let $\tau\geq 1$ be fixed. Let $q=L_{\mathrm{max}}/\sqrt{\ell\mu}$. Then PEARL-SGD with $\gamma_{k}\equiv\gamma=\frac{1}{\mu\eta(1+2q)}$ exhibits the rate
@@
Note that the $\tilde{{\mathcal{O}}}\left(\frac{(1+q)^{2}\left\|{\mathbf{x}}_{0}-{\mathbf{x}}_{\star}\right\|^{2}}{T^{2}}\right)$ term decays fast in Corollary 3.5 (as $T$ grows) and the terms proportional to $\sigma^{2}$ become dominant. The order of convergence is not slower than the $\tilde{{\mathcal{O}}}\left(\nicefrac{{1}}{{T}}\right)$ rate of the fully communicating case $\tau=1$, provided that $\nicefrac{{\tau^{2}L_{\mathrm{max}}\sigma^{2}}}{{\mu^{3}T^{2}}}={\mathcal{O}}\left(\nicefrac{{\sigma^{2}}}{{\mu^{2}T}}\right)\iff\tau={\mathcal{O}}\left(\sqrt{\nicefrac{{\mu T}}{{L_{\mathrm{max}}}}}\right)$. Therefore, as long as we select $\tau={\mathcal{O}}\left(\sqrt{\nicefrac{{\mu T}}{{L_{\mathrm{max}}}}}\right),$ in PEARL-SGD the communication cost is reduced by the factor of $\tau$ (because the total number of communications is $T/\tau$). With the largest possible $\tau$, the resulting communication complexity is $\nicefrac{{T}}{{\tau}}=\Theta\left(\sqrt{\nicefrac{{TL_{\mathrm{max}}}}{{\mu}}}\right)=\Theta\left(\sqrt{T}\right)$.
@@
We conclude the section with convergence result for PEARL-SGD using a decreasing step-size selection. While showing a similar convergence rate in terms of $T$ as in Corollary 3.5, Theorem 3.6 has the advantage of not requiring to fix $T$ in advance to determine the step-sizes.
@@
Under the assumptions of Theorem 3.4, let $q=L_{\mathrm{max}}/\sqrt{\ell\mu}$, and choose the step-sizes $\gamma_{k}=\begin{cases}\frac{1}{\ell\tau(1+2q)}&\text{if }p<2(1+2q)\kappa\\ \frac{1}{\tau\mu}\frac{2p+1}{(p+1)^{2}}&\text{if }p\geq 2(1+2q)\kappa\end{cases}$ for $\tau p\leq k\leq\tau(p+1)-1$, $p=0,\dots,R-1$. Then PEARL-SGD converges with the rate

### v2

@@ 3 PEARL-SGD: Algorithm and Convergence Guarantees
In this section, we introduce and analyze Algorithm˜1, named Per-Player Local SGD (PEARL-SGD), which is suitable for the MpFL setting we described in Section 2.
@@ 3.1 Algorithm and assumptions
In PEARL-SGD, clients/players of the game run SGD independently in parallel to update their strategy (keeping the strategies $x^{-i}$ of the other players fixed), which are communicated to other players only once in a while. In more detail, in every round of PEARL-SGD, each player $i\in[n]$ runs $\tau$ iterations of stochastic gradient descent (SGD) with respect to $f_{i}(\cdot,x^{-i})$, having $x^{-i}$ fixed to be the information of the other players’ actions obtained from the previous synchronization step. Once each player completes $\tau$ iterations of SGD (local updates), a synchronization occurs; the central server collects the actions of all players, and then the concatenation of all updated strategies/actions is distributed to all clients/players.
@@
Input: Step-sizes $\gamma_{k}>0$, Synchronization interval $\tau\geq 1$, Number of rounds $R\geq 1$
@@
  Master collects $x_{\tau p}^{i}$ from players $i\in[n]$
  Master distributes ${\mathbf{x}}_{\tau p}$ back to players
@@
Let us note that the synchronization step in PEARL-SGD involves transferring a $D=(d_{1}+\dots+d_{n})$-dimensional vector ${\mathbf{x}}_{\tau p}$ from the master server to the players. This is an important difference compared to the classical FL (minimization problem), where the dimension of the communication vectors is the same from client to master and from master to client, and it does not scale with $n$. While PEARL-SGD aims to reduce this overhead compared to its distributed variant ($\tau=1$) by communicating less frequently (with $\tau>1$), the high complexity of the synchronization step makes MpFL (and distributed $n$-player games in general) more suitable for cross-silo FL setups with relatively small number of organizations and more reliable communication bandwidths. We expect that the potentially expensive communication of high-dimensional vectors ${\mathbf{x}}_{\tau p}$ could be addressed by incorporating additional techniques such as gradient compression [alistarh2017qsgd, BeznosikovGorbunovBerardLoizou2023_stochastic]. This is an orthogonal approach to our proposed local methods, and we leave it for future work.
@@
There exists ${\mathbf{x}}_{\star}=(x_{\star}^{1},\dots,x_{\star}^{n})\in\mathbb{R}^{D}$, an equilibrium where ${\varbbF}({\mathbf{x}}_{\star})=0$ and $\mu>0$ such that $\forall{\mathbf{x}}\in\mathbb{R}^{D}$, $\left\langle{\varbbF}({\mathbf{x}}),{\mathbf{x}}-{\mathbf{x}}_{\star}\right\rangle\geq\mu\left\|{\mathbf{x}}-{\mathbf{x}}_{\star}\right\|^{2}$.
(QSM) is a concept extending quasi-strong convexity [gower2019sgd, GowerSebbouhLoizou2021_sgd] to the context of variational inequality problems (VIPs). This condition has been referred to as different names in the literature, such as strong coherent VIPs [song2020optimistic], VIPs with strong stability condition [mertikopoulos2019learning], or the strong Minty variational inequality [diakonikolas2021efficient]. It generalizes strong monotonicity, capturing some non-monotone problems. In [LoizouBerardGidelMitliagkasLacoste-Julien2021_stochastic], it was proposed and utilized as an assumption ensuring the convergence of SGDA dynamics in minimax games without the well-known issues of cycling or diverging [MeschederNowozinGeiger2017_numerics, DaskalakisIlyasSyrgkanisZeng2018_training]. Later, it was also used in the analysis of stochastic extragradient method [gorbunov2022stochastic] and its single-call variants (optimistic and past stochastic etragradient) [choudhury2024single].
@@
(SCO) generalizes the class of coercive operators and, interestingly, can hold for non-Lipschitz operators [LoizouBerardGidelMitliagkasLacoste-Julien2021_stochastic]. This has also been taken as minimal assumption for SGDA analysis in prior work [BeznosikovGorbunovBerardLoizou2023_stochastic]. Note that (QSM) and (SCO) together imply $\mu\left\|{\mathbf{x}}-{\mathbf{x}}_{\star}\right\|\leq\left\|{\varbbF}({\mathbf{x}})\right\|\leq\ell\left\|{\mathbf{x}}-{\mathbf{x}}_{\star}\right\|$ for any ${\mathbf{x}}\in\mathbb{R}^{D}$, which implies $\mu\leq\ell$. We call $\kappa=\ell/\mu\geq 1$ the condition number of the problem. In Appendix˜F, we provide a detailed discussion on the set of our theoretical assumptions and explain connections to other commonly assumed properties in the literature such as cocoercivity, Lipschitzness and monotonicity.
@@ 3.2 Convergence of PEARL-SGD: Deterministic setup
First, we provide the convergence result for PEARL-SGD with constant step-size $\gamma_{k}\equiv\gamma$ in the full-batch (deterministic) scenario, where there is no noise in the gradient computation. While this is directly recovered as a special case of Theorem˜3.4 with $\sigma_{i}=0$, we state it separately as the deterministic case provides several points of discussion that are worth emphasizing on their own.
@@
Assume (CVX), (SM), (QSM) and (SCO). Let $L_{\mathrm{max}}=\max\{L_{1},\dots,L_{n}\}$, $\kappa=\ell/\mu$ and $0<\gamma_{k}\equiv\gamma\leq\frac{1}{\ell\tau+2(\tau-1)L_{\mathrm{max}}\sqrt{\kappa}}$. Then the Deterministic PEARL-SGD (Algorithm˜1 with full-batch) converges with the rate
@@
Theorem 3.3 shows that deterministic PEARL-SGD converges linearly to an equilibrium. This is in contrast with the case of local gradient descent in FL setup with heterogeneous data, where convergence is only to a neighborhood of the optimum even in the absence of noise [KhaledMishchenkoRichtarik2020_tighter], unless further correction mechanism is used [MishchenkoMalinovskyStichRichtarik2022_proxskip]. This is because the classical FL problem is modeled as finite sum minimization, whereas our MpFL is modeled as a game, for which the existence of a variationally stable equilibrium is a standard assumption for convergence analysis [bravo2018bandit, mertikopoulos2019learning, LoizouBerardGidelMitliagkasLacoste-Julien2021_stochastic, mertikopoulos2024unified]. In particular, when $\tau=1$, the step-size constraint and the convergence rate of Theorem 3.3 reduces to the analysis of gradient descent-ascent (GDA) under the (QSM) and (SCO) assumptions from [LoizouBerardGidelMitliagkasLacoste-Julien2021_stochastic], showing that our analysis is tight and consistent with the existing literature.
@@
If $\gamma$ does not appropriately scale down with $\tau$, then at each round, players’ actions (SGD iterates) converge to minimizers of local functions. We call this phenomenon player drift, analogous to client drift in classical FL [KarimireddyKaleMohriReddiStichSuresh2020_scaffold], enforcing the ${\mathcal{O}}(1/\tau)$ step-size. In our setting, note that the local minimizers $x^{i}_{\star}(x_{\tau p}^{-i}):=\operatorname*{argmin}_{x^{i}\in\mathbb{R}^{d_{i}}}f_{i}(x^{i};x_{\tau p}^{-i})$ depend on other players’ strategies $x_{\tau p}^{-i}$. Due to this dependence, under extreme player drift, PEARL-SGD may display undesirable dynamics such as diverging to infinity (which can be checked with simple examples such as the two-player quadratic minimax game $\min_{u\in\mathbb{R}}\max_{v\in\mathbb{R}}\,\frac{\mu}{2}u^{2}+uv-\frac{\mu}{2}v^{2}$ with $\mu<1$). As these features are not typically observed in client drift in classical FL, player drift represents a distinct phenomenon despite some conceptual similarities. Therefore, we consider understanding and mitigating player drift an intriguing direction for future work in MpFL, which may necessitate novel insights that differ from existing approaches to client drift [KarimireddyKaleMohriReddiStichSuresh2020_scaffold, MishchenkoMalinovskyStichRichtarik2022_proxskip].
With the step-size constrained to $\gamma={\mathcal{O}}(1/\tau)$, communication reduction by PEARL-SGD is not observed in the deterministic setting (Theorem˜3.3) but is achieved in the stochastic setting—see Section˜3.3. A concurrent work [zindari2025decoupled] proposes the Decoupled SGD algorithm, which coincides with our PEARL-SGD, and shows that it can be communication-efficient even in the deterministic setup under an additional assumption of weakly coupled games (with slightly different main assumptions). We provide a more complete discussion of this in Appendix˜A.
@@ 3.3 Convergence of PEARL-SGD: Stochastic setup
We now discuss the convergence of PEARL-SGD with stochastic gradients. We first present the convergence of PEARL-SGD to a neighborhood of the equilibrium $\mathbf{x}$$_{\star}$ given constant step-sizes $\gamma_{k}\equiv\gamma$, and then discuss the communication complexity gain we achieve. Then we present the convergence result using a decreasing step-size selection, showing exact sublinear convergence to $\mathbf{x}$$_{\star}$ (rather than its neighborhood). We provide the outline and details of the proofs in Appendix˜C.
@@
If we use the largest possible step-size $\gamma=\frac{1}{\ell\tau+2(\tau-1)L_{\mathrm{max}}\sqrt{\kappa}}$ allowed in Theorem˜3.4, then the right-hand side of the bound does not scale down indefinitely with $\tau$. In fact, with this choice of $\gamma$, one can expect the communication gain by a factor of approximately $\frac{L_{\mathrm{max}}}{\ell}$ (when $L_{\mathrm{max}}\ll\ell$). More precisely, suppose $q\leq 1$ (equivalently $L_{\mathrm{max}}\leq\sqrt{\ell\mu}$—refer to Appendix˜F for the explanation that this is a common parameter regime). Then we have $\gamma=\Theta\left(\frac{1}{\ell\tau}\right)$ and
@@
In Corollary˜3.5, we analyze the convergence and communication gain of PEARL-SGD in the regime where the total number of iterations $T=\tau R$ is large, using a step-size depending on $T$.
@@
Suppose the assumptions of Theorem˜3.4 hold, and let $\tau\geq 1$ be fixed. Let $q=L_{\mathrm{max}}/\sqrt{\ell\mu}$. Then PEARL-SGD with $\gamma_{k}\equiv\gamma=\frac{1}{\mu\eta(1+2q)}$ exhibits the rate
@@
Note that the $\tilde{{\mathcal{O}}}\left(\frac{(1+q)^{2}\left\|{\mathbf{x}}_{0}-{\mathbf{x}}_{\star}\right\|^{2}}{T^{2}}\right)$ term decays fast in Corollary˜3.5 (as $T$ grows) and the terms proportional to $\sigma^{2}$ become dominant. The order of convergence is not slower than the $\tilde{{\mathcal{O}}}\left(\nicefrac{{1}}{{T}}\right)$ rate of the fully communicating case $\tau=1$, provided that $\nicefrac{{\tau^{2}L_{\mathrm{max}}\sigma^{2}}}{{\mu^{3}T^{2}}}={\mathcal{O}}\left(\nicefrac{{\sigma^{2}}}{{\mu^{2}T}}\right)\iff\tau={\mathcal{O}}\left(\sqrt{\nicefrac{{\mu T}}{{L_{\mathrm{max}}}}}\right)$. Therefore, as long as we select $\tau={\mathcal{O}}\left(\sqrt{\nicefrac{{\mu T}}{{L_{\mathrm{max}}}}}\right),$ in PEARL-SGD the communication cost is reduced by the factor of $\tau$ (because the total number of communications is $T/\tau$). With the largest possible $\tau$, the resulting communication complexity is $\nicefrac{{T}}{{\tau}}=\Theta\left(\sqrt{\nicefrac{{TL_{\mathrm{max}}}}{{\mu}}}\right)=\Theta\left(\sqrt{T}\right)$.
@@
We conclude the section with convergence result for PEARL-SGD using a decreasing step-size selection. While showing a similar convergence rate in terms of $T$ as in Corollary˜3.5, Theorem˜3.6 has the advantage of not requiring to fix $T$ in advance to determine the step-sizes.
@@
Under the assumptions of Theorem˜3.4, let $q=L_{\mathrm{max}}/\sqrt{\ell\mu}$, and choose the step-sizes $\gamma_{k}=\begin{cases}\frac{1}{\ell\tau(1+2q)}&\text{if }p<2(1+2q)\kappa\\ \frac{1}{\tau\mu}\frac{2p+1}{(p+1)^{2}}&\text{if }p\geq 2(1+2q)\kappa\end{cases}$ for $\tau p\leq k\leq\tau(p+1)-1$, $p=0,\dots,R-1$. Then PEARL-SGD converges with the rate